OpenStudy (anonymous):

find the turning point of y=2x^2+4x+3 i knoiw that y=3

OpenStudy (bahrom7893):

plug in y=3, and solve for x

OpenStudy (anonymous):

ah so simple

OpenStudy (anonymous):

i really stuffed up can you tell me what the turning point is cos i dont know if the teacher made a mistake'

OpenStudy (bahrom7893):

hmm, i don't feel like solving this equation, so i'll use a shortcut which u'll learn in calculus 1: 4x+4=0 x=1

OpenStudy (bahrom7893):

wait y is not 3 btw, when x is 1: y = 2+4+3=9

OpenStudy (anonymous):

i got y by making x =0

OpenStudy (bahrom7893):

well the turning point is: (1,9)

OpenStudy (anonymous):

well if ur right the teacher is wong he recons it is -1,1

OpenStudy (bahrom7893):

can't be. Plot and see.

OpenStudy (anonymous):

i will

OpenStudy (bahrom7893):

wait a sec, ur teacher's right.

OpenStudy (anonymous):

oh h

OpenStudy (anonymous):

how

OpenStudy (bahrom7893):

i made a mistake: 4x+4=0, x=-1

OpenStudy (anonymous):

ti will try complete the square

OpenStudy (bahrom7893):

Up there I wrote: 4x+4=0 x=1, x is -1, not 1

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i did it on my cas calculator and it says that it is -1,1

OpenStudy (bahrom7893):

yes that's right

OpenStudy (anonymous):

cos i dont know calculus i will try and solve it some way else

OpenStudy (anonymous):

ey i just use x=-b/2a

OpenStudy (anonymous):

whatever i do i wlways get y=3

OpenStudy (anonymous):

i got x=-1 sub that in and it makes y=3

OpenStudy (bahrom7893):

no when x=-1 y is not 3. Let's see 2x^2+4x+3 2*1 - 4 + 3 = 5 - 4 = 1

OpenStudy (anonymous):

ur a geneous i was taking 2x^2 as me having to square the 2 aswell

OpenStudy (bahrom7893):

I'm not a genius (facepalm) lol

OpenStudy (anonymous):

dont be modest